Please help me with this equation, I am not sure how to solve it. The diameters of red delicious apples of an orchard have a normal distribution with a mean of 3 inches and a standard deviation of 0.5 inch. What diameter measurement separates the smallest 33% from the largest 67%?
Found 2 solutions by rothauserc, Rogue-1:
Answer by rothauserc(4718)

We are given that
:
Pr(X<3) = 0.33
:
z-value for (X - 3) / 0.5 is -0.44, then
:
-0.44 = (X - 3) / 0.5
:
X = 3 − 0.22 = 2.78
:
now get z-value for 0.67 = 0.44
:
0.44 = (X - 3) / 0.5
:
X = 3 + 0.22 = 3.22
:
***********************
3.22 - 2.78 = 0.44
***********************

Answer by Rogue-1(1)

The diameters of red delicious apples of an orchard have a normal distribution with a mean of 3 inches and a standard deviation of 0.5 inch. One apple will be randomly chosen. What is the probability of picking an apple with diameter less than 2.15 inches? show step-by-step to derive to answer
-1.7
0.9554
0.0446
-3.85